A 25 kg child is sitting on a 10 kg sled. The child’s older sister is pulling the sled horizontally across the snow at a constant speed of 4 m/s by pulling on a rope with a force of 25 N at an angle of 30o above the horizontal. 

  

[4]   a.   Draw all appropriate diagrams for the situation. 

  

[3]   b.   Find the normal force acting on the sled. 

  

[3]   c.   Find the magnitude of the force of friction between the sled and the snow. 


1
Expert's answer
2022-02-22T08:53:36-0500


The normal reaction

"N=mg-F\sin 30^\circ\\
=25*9.8-25*0.5=232.5\:\rm N"

The friction force

"F_f=F\cos30^\circ=25*\sqrt{3}/2=21.7\:\rm N"